OpenStudy (anonymous):

find the circumferance of a circle whose area is 100m. please help :) thanks so much

OpenStudy (amistre64):

C = 2pi r A = pi r^2 100/pi = r^2 C = 2pi(sqrt(100/pi))

OpenStudy (amistre64):

20pi ------ looks wierd lol sqrt(pi)

OpenStudy (amistre64):

its abt 35.5

OpenStudy (anonymous):

its radius is 10 over square root of pi. so its circumference is 2pir 2pi(10/square root of pi) = 35.45

OpenStudy (anonymous):

pi r^2 = area of circle. pi r^2 = 100 Solve for r above.\[r=\frac{10}{\sqrt{\pi }} \]\[c=\pi\text{ 2 r}=\pi\text{ 2}\left(\frac{10}{\sqrt{\pi }}\right)=20 \sqrt{\pi } \]